What is the mass, in grams, of 1.50 mol of iron (II) sulfate? Express your answer using three significant figures m- 312 Submit My Answers Give Up Incorrect; Try Again; 5 attempts remaining Part B How many moles of ammonium ions are in 6935 g of ammonium carbonate? Express your answer using four significant figures. 0回 mol Submit My Answers Give Up Part C What is the mass, in grams, of 1.40-101 molecules of aspirin, CaH0,? Express your answer using three significant figures. Submit My Answers Give Up
Answer #1

A. Mass = mole * molar mass

= 1.5 * 400

= 600 or 6.00*10^2 grams

B. Moles of ammonium carbonate = mass / molar mass

= 6.935 / 96

= 0.0722

(NH4)2CO3 = 2NH4+ + CO32-

Moles of ammonium ion = 2* 0.072 = 0.144

C. Moles of aspirin = number of molecules / avagadro number

= 1.40*10^21 / (6.022*10^23)

= 2.325*10^-3

Mass of aspirin = mole * molar mass

= 2.325*10^-3 * 180.158

= 0.419 grams

If you have any query please comment

If you satisfied with the solution please like it thankxx